OpenStudy (anonymous):

Please Help me! Please! What is the inverse of the given relation? y = 3x + 9

OpenStudy (gorv):

ok now solve with me r u there???

OpenStudy (gorv):

@msm104

OpenStudy (anonymous):

yes im here

OpenStudy (gorv):

for calcuting inverse of this first calculate value of y in gterms of x

OpenStudy (gorv):

hey if u will calculate it will help u

OpenStudy (gorv):

find x= wat in terms of y

OpenStudy (gorv):

for finding x move each term on y side but not x

OpenStudy (gorv):

y=3x+9 y-9=3x (y-9)/3=x now for inverse replace x by y (x-9)/3=y that is your inverse

OpenStudy (anonymous):

A. \[y=\frac{ 1 }{ 3 }x -3\]

OpenStudy (anonymous):

B. \[y=3x-3\]

OpenStudy (anonymous):

C. \[y=3x+3\]

OpenStudy (gorv):

(x-9)/3=x/3-3 so A is the ans

OpenStudy (anonymous):

D. \[y= \frac{ 1 }{ 3 }x -3\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (gorv):

welcome buddy u got it???

OpenStudy (anonymous):

Yes sir